MILLIMETER WAVE SPECTROSCOPY OF THE INTERNAL ROTATION BANDS OF He-HCN AND He-DCN.

\underline{KENSUKE HARADA}, KEIICHI TANAKA, and TAKEHIKO TANAKA, Department of Chemistry, Faculty of Sciences, Kyushu University, Hakozaki, Higashiku, Fukuoka, 812-8581 JAPAN; SHINKO NANBU AND MUTSUMI AOYAGI, Institute for Molecular Science, Okazaki, 444-8585, Japan.

Millimeter-wave absorption spectroscopy combined with a pulsed-jet expansion technique was applied to the measurement of the internal-rotation band of the He-HCN complex in the frequency region of 95–125 GHz. In total 13 rovibrational lines, split into nitrogen nuclear hyperfine structure, were observed for the fundamental internal-rotation band, j = 1 − 0. The observed transition frequencies and the previous MBER data^{a} were analyzed to yield an empirical intermolecular potential energy surface, which was obtained by improving the recent ab initio potential surface.^{b} The surface obtained has a global minimum in the linear configuration (He⋯H–C–N) with a well depth of 30.2 cm^{−1}, and the saddle point located in the anti-linear configuration (H–C–N⋯He) is higher in energy than the global minimum by 8.92 cm^{−1}. The distance R_{m} from the He atom to the center of mass of HCN along the minimum energy path shows a large angular dependence; R_{m} is 4.169 Å and 4.039 Å in the linear and anti-linear forms, respectively, and has a minimum value of 3.528 Å in a T-shaped configuration. The He-DCN was also measured and analyzed. The empirical potential energy surfaces are compaired with the ab initio potential energy surface^{b} for both isotopic species.

aS. Drucker, F. Tao, and W. Klemperer, J. Phys. Chem. 99, 2646(1995).

bR. R. Toczylowski, F. Doloresco, and S. M. Cybulski, J. Chem. Phys. 114, 851(2001).